# A Posteriori Single- and Multi-Goal Error Control and Adaptivity for Partial Differential Equations

### v2

@@ 3.7 Error Estimation using Interpolation Techniques
~$\begin{split}\tilde{\mathcal{R}}^{(3)(2)}:=\frac{1}{2}\int_{0}^{1}[&J_{h}^{\prime\prime\prime}(\tilde{u}+se_{I_{h,u}})(e_{I_{h,u}},e_{I_{h,u}},e_{I_{h,u}})\\ -&(\mathcal{A}^{(2)})^{\prime\prime\prime}(\tilde{u}+se_{I_{h,u}})(e_{I_{h,u}},e_{I_{h,u}},e_{I_{h,u}},\tilde{z}+se_{I_{h,z}}^{*})\\ -&3(\mathcal{A}^{(2)})^{\prime\prime}(\tilde{u}+se_{I_{h,u}})(e_{I_{h,u}},e_{I_{h,u}},e_{I_{h,u}})]s(s-1)\,ds,\end{split}$
@@ 3.8.5 The Iteration Error Estimator $\eta_{I,h}$ for Interpolation
$\frac{1}{1+b_{h}^{I}}\left(|\eta_{I,h}|-|\eta_{k}+\eta_{\mathcal{R}_{I}}+\eta_{I_{u}}+\eta_{I_{z}}|\right)\leq|J(u)-J(\tilde{u})|\leq\frac{1}{1+b_{h}^{I}}\left(|\eta_{I,h}|+|\eta_{k}+\eta_{\mathcal{R}_{I}}+\eta_{I_{u}}+\eta_{I_{z}}|\right).$
@@
with $\underline{c}_{h}:=(1-\alpha_{\eta_{I,h}})\frac{1}{1+b_{h}^{I}}$ and $\overline{c}_{h}:=(1+\alpha_{\eta_{I,h}})\frac{1}{1+b_{h}^{I}}$.
@@ 3.10.2 Details on Using $\tilde{u}_{h}^{(2)}$ for the Localization
Remark 3.
In theory, there are no specific assumptions on the mesh except the usual requirements in terms of the (adaptive) finite element method like shape-regularity of the mesh; see Section 2.2. However, to make the discretization fully computable, so-called variational crimes are permitted; see, e.g., Chapter 8 in [55], For instance, one approximates the boundary $\partial\Omega$ of the spatial domain $\Omega$ or/and interfaces appropriately, and one uses numerical integration for computing the stiffness matrices, the right-hand sides, and the finite element approximation to the non-linear operator $\mathcal{A}(\cdot)$ and its computation at a finite element function / vector. This in turn depends on the mesh, and leads to additional contributions to the oscillation error. Furthermore, we have to assume that this error is sufficiently small such that the higher-order parts of the error decay fast. However, on very coarse meshes, this part can influence the discretization error estimator. We also refer to Section 4 for the analysis of non-standard discretizations.
@@ 4.2 Estimating the Consistency Error
$\rho_{h}(u)(v)=-\mathcal{A}_{h}(u)(v),\quad\rho_{h}^{*}(u,z)(v)=J^{\prime}(u)(v)-\mathcal{A}^{\prime}(u)(v,z).$
@@ 4.2.1 Consistency Error in Galerkin Time-Stepping Methods
$\mathcal{A}_{h}(u_{h})(v_{h})=\left(u_{h}(t_{0})-u_{0}\right)\cdot v_{h}(t_{0})+\sum_{n=1}^{N}\left(u_{h}(t_{n})-u_{h}(t_{n-1})+\Delta tf\left(t_{n},u(t_{n})\right)\right)\cdot v_{h}(t_{n}).$ (74)
@@
$\mathcal{A}(u_{h})(v_{h})=\left(u_{h}(t_{0})-u_{0}\right)\cdot v_{h}(t_{0})+\sum_{n=1}^{N}\left\{\left(u_{h}(t_{n})-u_{h}(t_{n-1}\right)\cdot v_{h}(t_{n})+\int_{t_{n-1}}^{t_{n}}f(t,u_{h}(t))\cdot v_{h}(t)\right\}.$ (76)
@@
The consistency error can be introduced as $\mathcal{S}_{h}(u)(z)=\mathcal{A}_{h}(u)(z)-\mathcal{A}(u)(z)$ and for the Euler method we obtain
@@ 4.3.1 Non-Conforming Finite Elements
4.3.2 Non-Conforming Boundary Approximation with Nitsche’s Method
As a typical example of a non-conforming discretization, we consider the problem: Find $u\in U:=H^{1}_{0}(\Omega)$ such that
$-\Delta u+u^{2}=f\text{ in }\Omega,$ (82)
which we approximate using Nitsche’s method [171] to weakly enforce the boundary conditions, i.e. finding $u_{h}\in U_{h}$ where $U_{h}\subset H^{1}(\Omega)$ but not necessarily $u_{h}=0$ on $\partial\Omega$ such that in the weak form, we have
$A_{h}(u_{h})(v_{h}):=(\nabla u_{h},\nabla v_{h})+(u_{h}^{2},v_{h})-\langle\partial_{n}u_{h},v_{h}\rangle_{\partial\Omega}+\frac{\gamma}{h}\langle u_{h},v_{h}\rangle_{\partial\Omega}-(f,v_{h})\stackrel{{\scriptstyle!}}{{=}}0\quad\forall v_{h}\in U_{h}(\Omega),$ (83)
where $\gamma>0$ is a parameter controlling the penalization of $u_{h}=0$. The linearized adjoint to (82) is: Find $z\in U:=H^{1}_{0}(\Omega)$ such that
$-\Delta z+2uz=j^{\prime}(u)\text{ in }\Omega,$ (84)
where $j^{\prime}(u)$ is the strong form of the goal functional, i.e. $J^{\prime}(u)(v)=\langle j^{\prime}(u),v\rangle$, and which is itself discretized using Nitsche’s method in the discrete variational form
$A_{h}^{*}(u_{h})(v_{h},z_{h}):=(\nabla v_{h},\nabla z_{h})+(2u_{h}v_{h},z_{h})-\langle\partial_{n}z_{h},v_{h}\rangle_{\partial\Omega}+\frac{\gamma}{h}\langle z_{h},v_{h}\rangle_{\partial\Omega}\stackrel{{\scriptstyle!}}{{=}}J^{\prime}(u_{h})(v_{h}).$ (85)
We note that it does not hold $A^{\prime}_{h}(u)(v,z)=A^{*}_{h}(u)(v,z)$ as the primal Nitsche form (83) does not contain a symmetry term on the boundary. Here, discretizing with Nitsche and taking the adjoint does not commute.
Since $u=v=z=0$ on $\partial\Omega$ for $u,v,z\in U:=H^{1}_{0}(\Omega)$, the approach is a consistent method and the solutions $u,z\in U$ satisfy $A_{h}(u,v)=0$ and $A^{*}_{h}(u)(v,z)=J^{\prime}(u)(v)$ for $v\in U$. In the following, we will mostly work with the abstract formulations $A_{h}(\cdot)(\cdot)$ and make it clear when properties of the specific problem (82) are used. With the Lagrangian $\mathcal{L}_{h}(u,z)=J(u)-A_{h}(u)(z)$ we formulate the error identity following the usual pathway (see for instance the proof of Theorem 1)
$J(u)-J(u_{h})=\frac{1}{2}\big{(}\mathcal{L}_{h}^{\prime}(u,z)(u-u_{h},z-z_{h})+\mathcal{L}_{h}^{\prime}(u_{h},z_{h})(u-u_{h},z-z_{h})\big{)}-A_{h}(u_{h})(z_{h})+{\cal R}^{(3)}.$ (86)
The very first term however will not vanish due to the non-conformity as $u_{h}$ and $z_{h}$ are not necessarily zero on the complete boundary $\partial\Omega$. As $L^{\prime}_{h}(u,z)(u,z)=0$ what remains is
$\mathcal{L}_{h}^{\prime}(u,z)(u-u_{h},z-z_{h})=-\mathcal{L}_{h}^{\prime}(u,z)(u_{h},z_{h}).$ (87)
For now, we leave this term and discuss it a little further down using problem (82) as an example. The second term or (86) is also not the standard weighted residual, as the adjoint variational form (85) is not the linearized adjoint of (83). Instead, we write
$\mathcal{L}^{\prime}_{h}(u_{h},z_{h})(u-u_{h},z-z_{h})$ $=-A_{h}(u_{h},z-z_{h})+J^{\prime}(u_{h})(u-u_{h})-A^{*}_{h}(u_{h})(u-u_{h},z_{h})$
$\quad+[A_{h}^{*}-A_{h}^{\prime}](u_{h})(u-u_{h},z_{h}),$ (88)
where $[A_{h}^{*}-A_{h}^{\prime}](u)(v,z):=A_{h}^{*}(u)(v,z)-A_{h}^{\prime}(u)(v,z)$ is the adjoint consistency error.
If we localize by introducing interpolations we must again take into account the non-conformity and finally reach the identity
$J(u)-J(u_{h})=\frac{1}{2}\rho_{h}(u_{h})(z-I_{h}z)+\frac{1}{2}\rho_{h}^{*}(u_{h},z_{h})(u-I_{h}u)+\rho_{h}(u_{h})(z_{h})+{\cal R}^{(3)}\\ -\frac{1}{2}\mathcal{L}^{\prime}_{h}(u,z)(u_{h},z_{h})+\frac{1}{2}[A_{h}^{*}-A_{h}^{\prime}](u_{h})(u-u_{h},z_{h})+\frac{1}{2}\mathcal{L}_{h}^{\prime}(u_{h},z_{h})(I_{h}u-u_{h},I_{h}z-z_{h})$ (89)
Here, $\rho_{h}(\cdot)(\cdot)$ is the usual residual and $\rho_{h}^{*}(u,z)(v)=J^{\prime}(u)(v)-A_{h}^{*}(u)(v,z)$. The first lines contains the usual terms: weighted primal and dual residuals, the iteration error and, finally, the third-order remainder. The second line consists of three parts: a term measuring the non-conformity of Nitsche’s method, a separate adjoint non-conformity term and, finally, another iteration error. In comparison to $\rho_{h}(u_{h})(z_{h})$, the term
$\mathcal{L}^{\prime}_{h}(u_{h},z_{h})(I_{h}u-u_{h},I_{h}z-z_{h})=J^{\prime}(u_{h})(I_{h}u-u_{h},z_{h})-A_{h}^{\prime}(u_{h})(I_{h}u-u_{h},z_{h})-A_{h}(u_{h})(I_{h}z-z_{h})$ (90)
is negligible as it is of higher order. It is an iteration error, however, it is additionally weighted with the errors $I_{h}u-u_{h}$ and $I_{h}z-z_{h}$.
Corollary 2 (Error estimation for Nitsche’s method applied to the model problem).
Let $u,z\in U:=H^{1}_{0}(\Omega)$ be solutions to (82) and (84), $u_{h},z_{h}\in U_{h}\subset H^{1}(\Omega)$ be approximations given by Nitsche’s method (83) and (85). Then, it holds
$J(u)-J(u_{h})\approx\frac{1}{2}\rho_{h}(u_{h})(z-I_{h}z)+\frac{1}{2}\rho_{h}^{*}(u_{h},z_{h})(u-I_{h}u)+\rho_{h}(u_{h})(z_{h})+\langle\partial_{n}z_{h},u_{h}\rangle_{\partial\Omega}.$ (91)
Proof.
We omit the third order remainder and discuss the three terms in the second line of (89). For the first non-conformity term we get with integration by parts and using $u=z=0$ on $\partial\Omega$ that
$\mathcal{L}^{\prime}_{h}(u,z)(u_{h},z_{h})$ $=J^{\prime}(u)(u_{h})-(\nabla z,\nabla u_{h})-(2uz,u_{h})+\langle u_{h},z\rangle_{\partial\Omega}-\frac{\gamma}{h}\langle u_{h},z\rangle_{\partial\Omega}$ (92)
$\quad+(f,z_{h})-(\nabla u,\nabla z_{h})-(u^{2},z_{h})+\langle\partial_{n}u,z_{h}\rangle_{\partial\Omega}-\frac{\gamma}{h}\langle u,z_{h}\rangle_{\partial\Omega}$
$=J^{\prime}(u)(u_{h})-(-\Delta z+2uz,u_{h})-\langle\partial_{n}z,u_{h}\rangle_{\partial\Omega}$
$\quad+(f,z_{h})-(-\Delta u+u^{2},z_{h})-\langle\partial_{n}u,z_{h}\rangle_{\partial\Omega}+\langle\partial_{n}u,z_{h}\rangle_{\partial\Omega}=-\langle\partial_{n}z,u_{h}\rangle_{\partial\Omega}.$
As $z\in H^{1}_{0}(\Omega)$ is not available, we introduce ±$\pm z_{h}$ and neglect the remainder term $\langle\partial_{n}(z-z_{h}),u_{h}\rangle_{\partial\Omega}$ as it carries an additional order of convergence over the remaining term $\langle\partial_{n}z_{h},u_{h}\rangle_{\partial\Omega}$.
For the second last term of (89) it holds using $u=z=0$ on $\partial\Omega$ that
$[A_{h}^{*}-A_{h}^{\prime}](u_{h})(u-u_{h},z_{h})$ $=-\langle\partial_{n}z_{h},u-u_{h}\rangle_{\partial\Omega}+\langle\partial_{n}(u-u_{h}),z_{h}\rangle_{\partial\Omega}$
$=\langle\partial_{n}z_{h},u_{h}\rangle_{\partial\Omega}+\langle\partial_{n}(u-u_{h}),z_{h}\rangle_{\partial\Omega}.$ (93)
We will neglect the second term, as it is qualitatively the same as (92) but carries the additional error $\partial_{n}(u-u_{h})$. Combining the above gives the postulated error approximation (91) ∎
The last term in (91) can be localized and attributed to the error indicators, e.g. by using the partition of unity method from Section 3.10. An alternative interpretation of this term however is to measure the violation of the boundary condition $u_{h}=0$ and it can be taken in its entirety to control the penalization parameter $\gamma>0$.
@@ 5.3 Error Localization
$J_{\mathfrak{E}}(u)-J_{\mathfrak{E}}(\tilde{u})\approx\eta^{(2)}=\sum_{i=1}^{M}\frac{1}{2}\rho(\tilde{u})((z_{h}^{(2)}-\tilde{z})\chi_{i})+\frac{1}{2}\rho^{*}(\tilde{u},\tilde{z})((u_{h}^{(2)}-\tilde{u})\chi_{i})+\rho(\tilde{u})(\tilde{z}),$
@@ 6.1.4 Goal Functional (Quantity of Interest) and Adjoint Problem
and was obtained on a globally refined super-mesh. In this simple example, with the ansatz $u(x,y)=\sin(\pi x)\sin(\pi y)$, we could have constructed a corresponding right hand side $f$ and with that we would have obtained an analytical reference value. Computing reference values on super-refined meshes is usually the method when no analytical or manufactured solutions are available.
@@ 6.4.2 Space-Time Finite Element Discretization and Non-Linear Solver
We implemented the space-time finite element method using the finite element library MFEM [7]. We use a $P_{1}$ conforming ansatz space for the 2D+1 ($d=2$) and 3D+1 ($d=3$) case, i.e. we use a simplicial decomposition $\mathcal{T}_{h}$ of the $(d+1)$-dimensional space-time cylinder $Q\subset\mathbb{R}^{d+1}$. For $d=2$, i.e. $Q\subset\mathbb{R}^{3}$, we use the bisection algorithm by Arnold, Mukherjee and Pouly [11], while for $d=3$, i.e. $Q\subset\mathbb{R}^{4}$, we use the newest vertex bisection from Stevenson’s paper [212]. Both algorithms ensure the shape regularity of the generated sequence of meshes.
@@ 7 Conclusions and Outlook
In this chapter, we reviewed single- and multigoal-oriented error control and adaptivity. Our notation was kept in general terms such that stationary and non-stationary (space-time) situations are covered. First, we explained in detail single goal-oriented error estimation with the help of the dual-weighted residual method. The error estimators cover both discretization and non-linear iteration errors. Therein, some new theoretical results were presented as well. Prior efficiency and reliability results only require one saturation assumption, rather a strengthened condition. Next, we considered non-standard discretizations such as stabilization terms and non-consistencies (e.g., classical finite difference based time-stepping schemes) or non-conformal methods (e.g., finite differences or neural network approximations) of numerical schemes in the frame of goal-oriented error estimation. Third, we concentrated on multigoal-oriented error estimation in which we have put a lot of efforts in the last eight years due to advances and increasing interest of solving multiphysics partial differential equations and coupled variational inequality systems. Besides theoretical results, we provided several adaptive algorithms for single and multiple goal functional evaluations. In order to substantiate our developments, four numerical examples were designed and computationally analyzed. In the first example, namely Poisson’s problem, the implementation is fully open-source and follows the classical structure of deal.II tutorial steps. Therefore, this example has a flavor of educational purpose. In the last example, recent work on the space-time PU-DWR method is further extended to multiple-goal functionals and singularities in the solution.
Ongoing and future work is concerned with the extension to cover more terms in the error estimators that besides balancing discretization and non-linear iteration errors, as well linear iteration errors, model errors, and model order reduction techniques are covered. Such error estimators could be theoretically investigated again with the help of saturation assumptions. The development and the analysis of fast iterative solvers for the linear and non-linear systems of finite element equations arising in the single- and multi-goal DWR AFEM considered in this chapter are also interesting topics of further research work, in particular, for the space-time AFEM presented in Subsection 6.4. It is very natural to use a nested iteration framework that reuses the generated sequence of adaptive meshes for the solvers like in the full linear and non-linear multigrid solvers; see, e.g., [124], or [53] for the full approximation scheme (FAS). Moreover, one should think about smart techniques between the use of expensive enriched spaces based on refined meshes or high-order elements and interpolation that keep the quality of enriched spaces with the efficiency of the interpolation technique. In conjunction with high-performance parallel computing these yield important components to continue to solve efficiently and with desired accuracies multiphysics problems with physics-based discretizations and fast physics-based numerical solvers. Two such applications are space-time fluid-structure interaction and space-time dynamic phase-field fracture problems. Partial results are available to date, but full space-time adaptivity not yet. Both can serve as full order models within model order reduction, where the reduced order models are designed themselves with the help of goal-oriented error estimation. Such by far cheaper reduced models enhance significantly the efficiency within optimization loops such as optimal control, inverse design and parameter estimation.

### v3

@@ 3.7 Error Estimation using Interpolation Techniques
~$\begin{split}\tilde{\mathcal{R}}^{(3)(2)}:=\frac{1}{2}\int_{0}^{1}[&J_{h}^{\prime\prime\prime}(\tilde{u}+se_{I_{h,u}})(e_{I_{h,u}},e_{I_{h,u}},e_{I_{h,u}})\\ -&(\mathcal{A}^{(2)})^{\prime\prime\prime}(\tilde{u}+se_{I_{h,u}})(e_{I_{h,u}},e_{I_{h,u}},e_{I_{h,u}},\tilde{z}+se_{I_{h,z}}^{*})\\ -&3(\mathcal{A}^{(2)})^{\prime\prime}(\tilde{u}+se_{I_{h,u}})(e_{I_{h,u}},e_{I_{h,u}},e_{I_{h,z}}^{*})]s(s-1)\,ds,\end{split}$
@@ 3.8.5 The Iteration Error Estimator $\eta_{I,h}$ for Interpolation
$\frac{1}{1+b_{h}^{I}}\left(|\eta_{I,h}|-|\eta_{k}+\eta_{\mathcal{R}_{I}}+\eta_{I_{u}}+\eta_{I_{z}}|\right)\leq|J(u)-J(\tilde{u})|\leq\frac{1}{1-b_{h}^{I}}\left(|\eta_{I,h}|+|\eta_{k}+\eta_{\mathcal{R}_{I}}+\eta_{I_{u}}+\eta_{I_{z}}|\right).$
@@
with $\underline{c}_{h}:=(1-\alpha_{\eta_{I,h}})\frac{1}{1+b_{h}^{I}}$ and $\overline{c}_{h}:=(1+\alpha_{\eta_{I,h}})\frac{1}{1-b_{h}^{I}}$.
@@ 4.2 Estimating the Consistency Error
$\rho_{h}(u)(v)=-A_{h}(u)(v),\quad\rho_{h}^{*}(u,z)(v)=J^{\prime}(u)(v)-A^{\prime}(u)(v,z).$
@@ 4.2.1 Consistency Error in Galerkin Time-Stepping Methods
$A_{h}(u_{h})(v_{h})=\left(u_{h}(t_{0})-u_{0}\right)\cdot v_{h}(t_{0})+\sum_{n=1}^{N}\left(u_{h}(t_{n})-u_{h}(t_{n-1})+\Delta tf\left(t_{n},u(t_{n})\right)\right)\cdot v_{h}(t_{n}).$ (74)
@@
$A(u_{h})(v_{h})=\left(u_{h}(t_{0})-u_{0}\right)\cdot v_{h}(t_{0})\\ +\sum_{n=1}^{N}\left\{\left(u_{h}(t_{n})-u_{h}(t_{n-1}\right)\cdot v_{h}(t_{n})+\int_{t_{n-1}}^{t_{n}}f(t,u_{h}(t))\cdot v_{h}(t)\right\}.$ (76)
@@
The consistency error can be introduced as $\mathcal{S}_{h}(u)(z)=A_{h}(u)(z)-A(u)(z)$ and for the Euler method we obtain
@@ 5.3 Error Localization
$J_{\mathfrak{E}}(u)-J_{\mathfrak{E}}(\tilde{u})\approx\eta^{(2)}=\sum_{i=1}^{M}\frac{1}{2}\rho(\tilde{u})((z_{h}^{(2)}-\tilde{z})\chi_{i})+\frac{1}{2}\rho^{*}(\tilde{u},\tilde{z})((u_{h}^{(2)}-\tilde{u})\chi_{i})+\rho(\tilde{u})(\tilde{z})$
@@ 6.1.4 Goal Functional (Quantity of Interest) and Adjoint Problem
and was obtained on a globally refined super-mesh.
@@ 6.4.2 Space-Time Finite Element Discretization and Non-Linear Solver
We implemented the space-time finite element method using the finite element library MFEM [7]. We use a $P_{1}$ conforming ansatz space for the 2D+1 ($d=2$) and 3D+1 ($d=3$) case, i.e. we use a simplicial decomposition $\mathcal{T}_{h}$ of the $(d+1)$-dimensional space-time cylinder $Q\subset\mathbb{R}^{d+1}$.
@@ 7 Conclusions and Outlook
In this work, we reviewed single- and multigoal-oriented error control and adaptivity. Our notation was kept in general terms such that stationary and non-stationary (space-time) situations are covered. First, we explained in detail single goal-oriented error estimation with the help of the dual-weighted residual method. The error estimators cover both discretization and non-linear iteration errors. Therein, some new theoretical results were presented as well. Prior efficiency and reliability results only require one saturation assumption, rather a strengthened condition. Next, we considered non-standard discretizations such as stabilization terms and non-consistencies (e.g., classical finite difference based time-stepping schemes) or non-conformal methods (e.g., finite differences or neural network approximations) of numerical schemes in the frame of goal-oriented error estimation. Third, we concentrated on multigoal-oriented error estimation in which we have put a lot of efforts in the last eight years due to advances and increasing interest of solving multiphysics partial differential equations and coupled variational inequality systems. Besides theoretical results, we provided several adaptive algorithms for single and multiple goal functional evaluations. In order to substantiate our developments, four numerical examples were designed and computationally analyzed. In the first example, namely Poisson’s problem, the implementation is fully open-source and follows the classical structure of deal.II tutorial steps. Therefore, this example has a flavor of educational purpose. In the last example, recent work on the space-time PU-DWR method is further extended to multiple-goal functionals and singularities in the solution. Ongoing and future work is concerned with the extension to cover more terms in the error estimators that besides discretization and non-linear iteration errors, as well linear iteration errors, model errors, and model order reduction techniques are covered. In conjunction with high-performance parallel computing these yield important components to continue to solve efficiently and with desired accuracies multiphysics problems with physics-based discretizations and fast physics-based numerical solvers.